Fertilizer with weed control costing $36 per 50-pound bag is to be mixed with a less expensive fertilizer costing $8 per 50-pound bag to make 14 bags of fertilizer that can be sold for $28 per bag. How many bags of cheaper fertilizer should be used?
Answer by stanbon(75887)

Fertilizer with weed control costing $36 per 50-pound bag is to be mixed with a less expensive fertilizer costing $8 per 50-pound bag to make 14 bags of fertilizer that can be sold for $28 per bag. How many bags of cheaper fertilizer should be used?
-----
Bag Eq:: t + e = 14 bags
Value Eq:36t+8e = 14*28
--------
Modify for elimination::
8t + 8e = 8*14
36t + 8e = 28*14
------------------------
28t = 20*14
---------------
t = 10 (# of $36 bags)
e = 14-10 = 4 (# of $8 bags)
--------------
Cheers,
Stan H.
-------------